OpenStudy (anonymous):

The cost for laminating varies directly with the number of pages laminated. A customer had 19 pages laminated and paid $16.15. How much will the customer pay to have 31 pages laminated? $

OpenStudy (anonymous):

@DanJS can u help me please?

OpenStudy (danjs):

hi

OpenStudy (danjs):

First sentence - varies directly - y = k*x

OpenStudy (danjs):

we need to find the constant of proportionality ,,, k using the point they gave you.

OpenStudy (danjs):

Cost = k*# pages

OpenStudy (anonymous):

so what is the answer?

OpenStudy (danjs):

16.15 = k*19 k = 0.85 so Cost = 0.85*(#pages)

OpenStudy (danjs):

The cost of 31 pages will be: Cost = 0.85*(31 pages) = 26.35

OpenStudy (anonymous):

thxs:)

OpenStudy (danjs):

varies directly: \[y=k*x\] varies indirectly: \[y = k*\frac{ 1 }{ x }\]